OpenStudy (anonymous):

A freshwater tank holds 400 gallons and evaporates 2 gallons per week. A salt water tank holds 451 gallons and evaporates 5 gallons per week. In how many weeks will the two tanks hold the same amount? 7 9 15 17

OpenStudy (hoblos):

400 - 2x = 451-5x x=17

OpenStudy (anonymous):

thanks